# Poincaré inequality and Quantitative De Giorgi method for hypoelliptic operators

### v2

@@ 3. Proof of the Poincaré inequality
We consider the three cylinders $Q^{+},Q^{-},Q^{0}\subset\Omega$ as in Figure 2, operators $\mathcal{A}$ and $\mathcal{B}$ that satisfy (H), and a subsolution $f$ of (1.1). Let $\varphi\in C_{c}^{\infty}({\mathbb{R}}^{N})$ be a non-negative function in the first $N$ variables $(x^{(\kappa)},\dots,x^{(0)})$ (excluding the time component) with compact support in any time-slice of $Q^{0}$ and such that $\fint_{Q^{0}}\varphi{\,\mathrm{d}}x=1$. Then
@@
(3.2) ${\mathcal{I}}^{\pm}$ =±$=\pm\fint_{Q^{0}}\Big{(}f(z_{\pm})-f(z_{0})\Big{)}\,\varphi(x_{0}){\,\mathrm{d}}z_{0}$
@@
±$\leq\underbrace{\pm\delta_{\pm}\fint_{Q^{0}}\int_{0}^{1}\left[\left(\mathcal{A}^{*}\mathcal{A}\right)f\right]\big{(}\Gamma_{\pm}(s)\big{)}\,\varphi(x_{0}){\,\mathrm{d}}s{\,\mathrm{d}}z_{0}}_{=:{\mathcal{I}}^{\pm}_{1}}$
@@
Note that the only difference between the two terms ${\mathcal{I}}^{+}$ and ${\mathcal{I}}^{-}$ is the role of $z_{0}$: in the former case it is the past variable, in the latter it is the future variable.
In the following computations to estimate $\mathcal{I}_{1}^{\pm}$ and $\mathcal{I}_{2}^{\pm}$, the local and the non-local case can be treated similarly, upon replacing the integration domain over $Q^{0}$ by ${\mathbb{R}}^{d_{0}}\times Q^{0}_{v}$ in the non-local case, where $Q^{0}_{v}\subset{\mathbb{R}}^{N+1-d_{0}}$ is the domain of the cylinder $Q^{0}$ for the variables $(x^{(1)},\dots,x^{(\kappa)},t)$, that is we slice out the $x^{(0)}$ variable.
We now integrate by parts the terms ${\mathcal{I}}^{\pm}_{1}$ after a change of variables $x_{0}\mapsto y:=\Phi^{s}_{\pm}(x_{0})$ for $s,t_{0}$ fixed ($\Phi^{s}_{\pm}$ is the affine map defined in (2.11), which is invertible for $s\not=0$):
${\mathcal{I}}_{1}^{\pm}$ =±$=\pm\delta_{\pm}\fint_{Q^{0}}\int_{0}^{1}\left[\left(\mathcal{A}^{*}\mathcal{A}\right)f\right]\big{(}\Gamma_{\pm}(s)\big{)}\,\varphi(x_{0}){\,\mathrm{d}}s{\,\mathrm{d}}z_{0}$
=±$=\pm\delta_{\pm}\fint_{Q^{0}}\int_{0}^{1}\left[\left(\mathcal{A}^{*}\mathcal{A}\right)f\right]\left(\Phi^{s}_{\pm}(x_{0}),st_{0}+(1-s)t_{\pm}\right)\,\varphi(x_{0}){\,\mathrm{d}}s{\,\mathrm{d}}x_{0}{\,\mathrm{d}}t_{0}$
=±$=\pm\frac{\delta_{\pm}}{|Q^{0}|}\int_{(\Phi^{s}_{\pm}\otimes\text{Id})(Q^{0})}\int_{0}^{1}\left[\left(\mathcal{A}^{*}\mathcal{A}\right)f\right]\left(y,st_{0}+(1-s)t_{\pm}\right)\,\varphi\left(\left(\Phi^{s}_{\pm}\right)^{-1}(y)\right)\frac{{\,\mathrm{d}}s{\,\mathrm{d}}y{\,\mathrm{d}}t_{0}}{\left|\det\mathfrak{A}^{s}\right|}$
=±$=\pm\frac{\delta_{\pm}}{|Q^{0}|}\int_{(\Phi^{s}_{\pm}\otimes\text{Id})(Q^{0})}\int_{0}^{1}\left[\mathcal{A}f\right]\left(y,st_{0}+(1-s)t_{\pm}\right)\,{\mathcal{A}}\left[\varphi\left(\left(\Phi^{s}_{\pm}\right)^{-1}(y)\right)\right]\frac{{\,\mathrm{d}}s{\,\mathrm{d}}y{\,\mathrm{d}}t_{0}}{\left|\det\mathfrak{A}^{s}\right|}$
=±$=\pm\frac{\delta_{\pm}}{|Q^{0}|}\int_{Q^{0}}\int_{0}^{1}\left[\mathcal{A}f\right]\left(\Gamma_{\pm}(s)\right)\,\left\{{\mathcal{A}}\left[\varphi\left(\left(\Phi^{s}_{\pm}\right)^{-1}(y)\right)\right]\right\}_{|y=\Phi^{s}_{\pm}(x_{0})}{\,\mathrm{d}}s{\,\mathrm{d}}x_{0}{\,\mathrm{d}}t_{0}.$
We then use the first bound in (1.13) on the operator $\mathcal{A}$ in assumption (H):
|$\left|\left\{{\mathcal{A}}\left[\varphi\left(\left(\Phi^{s}_{\pm}\right)^{-1}(y)\right)\right]\right\}_{|y=(\Phi^{s}_{\pm})^{-1}(x_{0})}\right|$ $\lesssim_{\varphi}\left\|\nabla_{(0)}\left(\Phi^{s}_{\pm}\right)^{-1}\right\|_{\infty}^{\beta}$
$\lesssim_{\varphi}\left(\sum_{i=0}^{\kappa}s^{-1-\alpha_{i}}\right)^{\beta}\lesssim_{\varphi}\left(\sum_{i=0}^{\kappa}s^{-\beta-\beta\alpha_{i}}\right)$
@@
${\mathcal{I}}_{1}^{\pm}\lesssim|\delta_{\pm}|\fint_{Q^{0}}\int_{0}^{1}\left|\left[\mathcal{A}f\right]\left(\Gamma_{\pm}(s)\right)\right|\left(\sum_{i=0}^{\kappa}s^{-\beta-\beta\alpha_{i}}\right){\,\mathrm{d}}s{\,\mathrm{d}}x_{0}{\,\mathrm{d}}t_{0}.$
@@
×$\hskip 227.62204pt\times\varphi\left(\left(\Phi^{s}_{\pm}\right)^{-1}(y)\right)\frac{{\,\mathrm{d}}s{\,\mathrm{d}}y{\,\mathrm{d}}t_{0}}{\left|\det\mathfrak{A}^{s}\right|}$
@@
×$\hskip 170.71652pt\times(-\Delta_{(0)})^{\frac{1-\beta}{2}}\left\{\varphi\left[\left(\Phi_{\pm}^{s}\right)^{-1}(y)\right]\right\}\frac{{\,\mathrm{d}}s{\,\mathrm{d}}y{\,\mathrm{d}}t_{0}}{\left|\det\mathfrak{A}^{s}\right|}$
$=\mp\frac{1}{|Q^{0}|}\sum_{i=0}^{\kappa}\int_{0}^{1}\int_{(\Phi^{s}_{\pm}\otimes\text{Id})(Q^{0})}s^{\alpha_{i}}\mathsf{m}_{\pm}^{(i)}\cdot\left(\nabla_{(0)}(-\Delta_{(0)})^{-\frac{1-\beta}{2}}f\right)\left(\Gamma_{\pm}(s)\right)$
×$\hskip 99.58464pt\times\left\{(-\Delta_{(0)})^{\frac{1-\beta}{2}}\left\{\varphi\left[\left(\Phi_{\pm}^{s}\right)^{-1}(y)\right]\right\}\right\}_{|y=(\Phi^{s}_{\pm})^{-1}(x_{0})}{\,\mathrm{d}}s{\,\mathrm{d}}x_{0}{\,\mathrm{d}}t_{0}.$
We then observe that
$\left|(-\Delta_{(0)})^{\frac{1-\beta}{2}}\left\{\varphi\left[\left(\Phi_{\pm}^{s}\right)^{-1}(y)\right]\right\}\right|\lesssim\left\|\nabla_{(0)}\left(\Phi^{s}_{\pm}\right)^{-1}\right\|_{\infty}^{1-\beta}\lesssim\left(\sum_{j=0}^{\kappa}s^{-1-\alpha_{j}}\right)^{1-\beta}$
@@
${\mathcal{I}}_{2}^{\pm}\leq\frac{1}{|Q^{0}|}\int_{0}^{1}\int_{Q^{0}}\left(\sum_{i,j=0}^{\kappa}s^{\alpha_{i}+\beta(1+\alpha_{j})-\alpha_{j}-1}\right)\left|\left(\nabla_{(0)}(-\Delta_{(0)})^{-\frac{1-\beta}{2}}f\right)\left(\Gamma_{\pm}(s)\right)\right|{\,\mathrm{d}}s{\,\mathrm{d}}x_{0}{\,\mathrm{d}}t_{0}.$
@@
(3.3) ${\mathcal{J}}$ $\lesssim\int_{Q^{\pm}}\int_{0}^{1}\int_{Q^{0}}\left(\sum_{i=0}^{\kappa}\frac{1}{s^{(\alpha_{i}+1)\beta}}\right)\left|\left({\mathcal{A}}f\right)(\Gamma_{\pm}(s))\right|{\,\mathrm{d}}z_{\pm}{\,\mathrm{d}}s{\,\mathrm{d}}z_{0}$
$+\int_{Q^{\pm}}\int_{0}^{1}\int_{Q^{0}}\left(\sum_{i,j=0}^{\kappa}s^{\alpha_{i}+\beta(1+\alpha_{j})-\alpha_{j}-1}\right)\left|\left(\nabla_{(0)}(-\Delta_{(0)})^{-\frac{1-\beta}{2}}f\right)\left(\Gamma_{\pm}(s)\right)\right|{\,\mathrm{d}}z_{\pm}{\,\mathrm{d}}s{\,\mathrm{d}}z_{0}.$
In order to balance optimally the singularities we then choose the $\alpha_{i}$’s all close to $(2+\beta)^{-1}-1\in(-1,0)$ pairwise distinct.
@@
where we have used first that the integral in $s$ avoids the singularity at 00, and second assumption (1.13) in (H). To control the part ${\mathcal{J}}^{\pm}_{1}$ we parametrise $\Gamma_{\pm}(s)$ by the $z_{\pm}$ coordinates, for $s\in[0,s_{0}]$ and $t_{0},t_{\pm}$ all fixed:
Since $s$ does not approach $1$, we can prove that this change of variables in not singular. The mappings $\Psi_{\pm}^{s}$ are determined by solving (2.11) for $x_{\pm}$ instead of $x_{0}$. It yields $\Psi_{\pm}^{s}x_{\pm}=\mathfrak{a}_{\pm}^{s}x_{\pm}+\mathfrak{b}_{\pm}^{s}$ with the matrix $\mathfrak{a}_{\pm}^{s}$ and vector $\mathfrak{b}_{\pm}^{s}$ given by
$\mathfrak{a}^{s}_{\pm}:=\left({\mathsf{T}}_{\pm}(s)-{\mathsf{W}}_{\pm}^{\delta}(s)\big{[}{\mathsf{W}}_{\pm}^{\delta}(1)\big{]}^{-1}{\mathsf{T}}_{\pm}(1)\right)\quad\mathfrak{b}^{s}_{\pm}:={\mathsf{W}}_{\pm}^{\delta}(s)\big{[}{\mathsf{W}}_{\pm}^{\delta}(1)\big{]}^{-1}x_{0}$
$\lesssim_{s_{0}}\int_{\tilde{\Omega}}\left|{\mathcal{A}}f(z)\right|{\,\mathrm{d}}z$
which concludes the proof. Note that in the local case $\tilde{\Omega}$ coincides with the solution domain $\Omega$, whereas in the non-local case, the domain for the first $d_{0}$ variables of $\tilde{\Omega}$ coincides with the whole space ${\mathbb{R}}^{d_{0}}$, so that $\tilde{\Omega}={\mathbb{R}}^{d_{0}}\times B_{R}\times\dots\times B_{R}\times(-5,0)$.
4. Proof of the Harnack inequality
In this section we briefly explain how to adapt the approach of [8] for deducing the intermediate value of De Giorgi from the Poincaré inequality we have proved. This approach was developed in the local case with one commutator and with a weaker version of the Poincaré inequality. A non-local extension, with one commutator, was developed in [13]. This approach can be described as a “trajectory viewpoint” on the De Giorgi theory, and it is fully quantitative. The first De Giorgi lemma writes in its simplest form

### v3

@@ 3. Proof of the Poincaré inequality
We consider the three cylinders $Q^{+},Q^{-},Q^{0}\subset\tilde{\Omega}$ as in Figure 2, operators $\mathcal{A}$ and $\mathcal{B}$ that satisfy (H), and a weak sub-solution $f$ of (1.1). Let $\varphi\in C_{c}^{\infty}({\mathbb{R}}^{N})$ be a non-negative function in the first $N$ variables $(x^{(\kappa)},\dots,x^{(0)})$ (excluding the time component) with compact support in any time-slice of $Q^{0}$ and such that $\fint_{Q^{0}}\varphi{\,\mathrm{d}}x=1$. Then
@@
(3.2) ${\mathcal{I}}^{\pm}$ $=\pm\fint_{Q^{0}}\Big{(}f(z_{\pm})-f(z_{0})\Big{)}\,\varphi(x_{0}){\,\mathrm{d}}z_{0}$
@@
$\leq\underbrace{\pm\delta_{\pm}\fint_{Q^{0}}\int_{0}^{1}\left[\left(\mathcal{A}^{*}\mathcal{A}\right)f\right]\big{(}\Gamma_{\pm}(s)\big{)}\,\varphi(x_{0}){\,\mathrm{d}}s{\,\mathrm{d}}z_{0}}_{=:{\mathcal{I}}^{\pm}_{1}}$
@@
where the above integrals are interpreted in the duality sense. The only difference between the two terms $\mathcal{I}$$^{+}$ and $\mathcal{I}$$^{-}$ is the role of $z_{0}$: in the former case it is the past variable, in the latter it is the future variable.
In the following computations to estimate $\mathcal{I}_{1}^{\pm}$ and $\mathcal{I}_{2}^{\pm}$, the local and the non-local case can be treated similarly, upon replacing the integration domain over $Q^{0}$ by $Q^{0}_{v}\times{\mathbb{R}}^{d_{0}}$ in the non-local case, where $Q^{0}_{v}\subset{\mathbb{R}}^{N+1-d_{0}}$ is the domain of the cylinder $Q^{0}$ for the variables $(x^{(1)},\dots,x^{(\kappa)},t)$, that is we slice out the $x^{(0)}$ variable.
We now integrate by parts the terms ${\mathcal{I}}^{\pm}_{1}$ after a change of variables $x_{0}\mapsto y:=\Phi^{s}_{\pm}(x_{0})$ for $s,t_{0}$ fixed $\Phi^{s}_{\pm}$ is the affine map defined in (2.11), which is invertible for $s\not=0$):
${\mathcal{I}}_{1}^{\pm}$ $=\pm\frac{\delta_{\pm}}{|Q^{0}|}\int_{Q^{0}}\int_{0}^{1}\left[\left(\mathcal{A}^{*}\mathcal{A}\right)f\right]\big{(}\Gamma_{\pm}(s)\big{)}\,\varphi(x_{0}){\,\mathrm{d}}s{\,\mathrm{d}}z_{0}$
$=\pm\frac{\delta_{\pm}}{|Q^{0}|}\int_{Q^{0}}\int_{0}^{1}\left[\left(\mathcal{A}^{*}\mathcal{A}\right)f\right]\left(\Phi^{s}_{\pm}(x_{0}),st_{0}+(1-s)t_{\pm}\right)\,\varphi(x_{0}){\,\mathrm{d}}s{\,\mathrm{d}}x_{0}{\,\mathrm{d}}t_{0}$
$=\pm\frac{\delta_{\pm}}{|Q^{0}|}\int_{(\Phi^{s}_{\pm}\otimes\text{Id})(Q^{0})}\int_{0}^{1}\left[\left(\mathcal{A}^{*}\mathcal{A}\right)f\right]\left(y,st_{0}+(1-s)t_{\pm}\right)\,\varphi\left(\left(\Phi^{s}_{\pm}\right)^{-1}(y)\right)\frac{{\,\mathrm{d}}s{\,\mathrm{d}}y{\,\mathrm{d}}t_{0}}{\left|\det\mathfrak{A}^{s}\right|}$
$=\pm\frac{\delta_{\pm}}{|Q^{0}|}\int_{(\Phi^{s}_{\pm}\otimes\text{Id})(Q^{0})}\int_{0}^{1}\left[\mathcal{A}f\right]\left(y,st_{0}+(1-s)t_{\pm}\right)\,{\mathcal{A}}\left[\varphi\left(\left(\Phi^{s}_{\pm}\right)^{-1}(y)\right)\right]\frac{{\,\mathrm{d}}s{\,\mathrm{d}}y{\,\mathrm{d}}t_{0}}{\left|\det\mathfrak{A}^{s}\right|}$
$=\pm\frac{\delta_{\pm}}{|Q^{0}|}\int_{Q^{0}}\int_{0}^{1}\left[\mathcal{A}f\right]\left(\Gamma_{\pm}(s)\right)\,\left\{{\mathcal{A}}\left[\varphi\left(\left(\Phi^{s}_{\pm}\right)^{-1}(y)\right)\right]\right\}_{|y=\Phi^{s}_{\pm}(x_{0})}{\,\mathrm{d}}s{\,\mathrm{d}}x_{0}{\,\mathrm{d}}t_{0}.$
We then use the first bound in (1.14) on the operator $\mathcal{A}$ in assumption (H): for $\epsilon\in(0,1-\beta)$
$\left|\left\{{\mathcal{A}}\left[\varphi\left(\left(\Phi^{s}_{\pm}\right)^{-1}(y)\right)\right]\right\}_{|y=(\Phi^{s}_{\pm})^{-1}(x_{0})}\right|$ $\lesssim_{\varphi}\left\|\nabla_{(0)}\left(\Phi^{s}_{\pm}\right)^{-1}\right\|_{\infty}^{\beta+\epsilon}$
$\lesssim_{\varphi}\left(\sum_{i=0}^{\kappa}s^{-1-\alpha_{i}}\right)^{\beta+\epsilon}\lesssim_{\varphi}\left(\sum_{i=0}^{\kappa}s^{-(\beta+\epsilon)(1+\alpha_{i})}\right)$
@@
${\mathcal{I}}_{1}^{\pm}\lesssim\frac{\delta_{\pm}}{|Q^{0}|}\int_{Q^{0}}\int_{0}^{1}\left|\left[\mathcal{A}f\right]\left(\Gamma_{\pm}(s)\right)\right|\left(\sum_{i=0}^{\kappa}s^{-(\beta+\epsilon)(1+\alpha_{i})}\right){\,\mathrm{d}}s{\,\mathrm{d}}x_{0}{\,\mathrm{d}}t_{0}.$
@@
$\hskip 227.62204pt\times\varphi\left(\left(\Phi^{s}_{\pm}\right)^{-1}(y)\right)\frac{{\,\mathrm{d}}s{\,\mathrm{d}}y{\,\mathrm{d}}t_{0}}{\left|\det\mathfrak{A}^{s}\right|}$
@@
$\hskip 170.71652pt\times(-\Delta_{(0)})^{\frac{1-\beta}{2}}\left\{\varphi\left[\left(\Phi_{\pm}^{s}\right)^{-1}(y)\right]\right\}\frac{{\,\mathrm{d}}s{\,\mathrm{d}}y{\,\mathrm{d}}t_{0}}{\left|\det\mathfrak{A}^{s}\right|}$
$=\mp\frac{1}{|Q^{0}|}\sum_{i=0}^{\kappa}\int_{0}^{1}\int_{{Q^{0}}}s^{\alpha_{i}}\mathsf{m}_{\pm}^{(i)}\cdot\left(\nabla_{(0)}(-\Delta_{(0)})^{-\frac{1-\beta}{2}}f\right)\left(\Gamma_{\pm}(s)\right)$
$\hskip 99.58464pt\times\left\{(-\Delta_{(0)})^{\frac{1-\beta}{2}}\left\{\varphi\left[\left(\Phi_{\pm}^{s}\right)^{-1}(y)\right]\right\}\right\}_{|y=(\Phi^{s}_{\pm})^{-1}(x_{0})}{\,\mathrm{d}}s{\,\mathrm{d}}x_{0}{\,\mathrm{d}}t_{0}.$
We then observe that for some $\epsilon\in(0,1)$
$\left|(-\Delta_{(0)})^{\frac{1-\beta}{2}}\left\{\varphi\left[\left(\Phi_{\pm}^{s}\right)^{-1}(y)\right]\right\}\right|\lesssim\left\|\nabla_{(0)}\left(\Phi^{s}_{\pm}\right)^{-1}\right\|_{\infty}^{1-\beta-\epsilon}\lesssim\left(\sum_{j=0}^{\kappa}s^{-1-\alpha_{j}}\right)^{1-\beta-\epsilon},$
@@
${\mathcal{I}}_{2}^{\pm}\leq\frac{1}{|Q^{0}|}\int_{0}^{1}\int_{Q^{0}}\left(\sum_{i,j=0}^{\kappa}s^{\alpha_{i}+(\beta+\epsilon)(1+\alpha_{j})-\alpha_{j}-1}\right)\left|\left(\nabla_{(0)}(-\Delta_{(0)})^{-\frac{1-\beta}{2}}f\right)\left(\Gamma_{\pm}(s)\right)\right|{\,\mathrm{d}}s{\,\mathrm{d}}x_{0}{\,\mathrm{d}}t_{0}.$
@@
(3.3) ${\mathcal{J}}$ $\lesssim\int_{Q^{\pm}}\int_{0}^{1}\int_{Q^{0}}\left(\sum_{i=0}^{\kappa}\frac{1}{s^{(\alpha_{i}+1)(\beta+\epsilon)}}\right)\left|\left({\mathcal{A}}f\right)(\Gamma_{\pm}(s))\right|{\,\mathrm{d}}z_{\pm}{\,\mathrm{d}}s{\,\mathrm{d}}z_{0}$
$+\int_{Q^{\pm}}\int_{0}^{1}\int_{Q^{0}}\left(\sum_{i,j=0}^{\kappa}s^{\alpha_{i}+(\beta+\epsilon)(1+\alpha_{j})-\alpha_{j}-1}\right)\left|\left(\nabla_{(0)}(-\Delta_{(0)})^{-\frac{1-\beta}{2}}f\right)\left(\Gamma_{\pm}(s)\right)\right|{\,\mathrm{d}}z_{\pm}{\,\mathrm{d}}s{\,\mathrm{d}}z_{0}.$
In order to balance optimally the singularities we then choose the $\alpha_{i}$’s all close to $(2+\beta+\epsilon)^{-1}-1\in(-1,0)$ and pairwise distinct.
@@
To control the part ${\mathcal{J}}^{\pm}_{1}$ we parametrise $\Gamma_{\pm}(s)$ by the $z_{\pm}$ coordinates, for $s\in[0,s_{0}]$ and $t_{0},t_{\pm}$ all fixed:
Since $s$ does not approach $1$, we can prove that this change of variables is not singular. The mappings $\Psi_{\pm}^{s}$ are determined by solving (2.11) for $x_{\pm}$ instead of $x_{0}$. It yields $\Psi_{\pm}^{s}x_{\pm}=\mathfrak{a}_{\pm}^{s}x_{\pm}+\mathfrak{b}_{\pm}^{s}$ with the matrix $\mathfrak{a}_{\pm}^{s}$ and vector $\mathfrak{b}_{\pm}^{s}$ given by
$\mathfrak{a}^{s}_{\pm}:=\left({\mathsf{T}}_{\pm}(s)-{\mathsf{W}}_{\pm}^{\delta}(s)\big{[}{\mathsf{W}}_{\pm}^{\delta}(1)\big{]}^{-1}{\mathsf{T}}_{\pm}(1)\right)\quad\mathfrak{b}^{s}_{\pm}:={\mathsf{W}}_{\pm}^{\delta}(s)\big{[}{\mathsf{W}}_{\pm}^{\delta}(1)\big{]}^{-1}x_{0},$
$\lesssim_{s_{0}}\int_{\tilde{\Omega}}\left|{\mathcal{A}}f(z)\right|{\,\mathrm{d}}z.$
@@
where we have used first that the integral in $s$ avoids the singularity at 0, and second assumption (1.14) in (H). This concludes the proof.
4. Proof of the weak Harnack inequality
In this section we briefly explain how to adapt the approach of [8] for deducing the intermediate value of lemma of De Giorgi from the Poincaré inequality we have proved. This approach was developed in the local case with $\kappa=1$ and with a weaker version of the Poincaré inequality. A non-local extension, when $\kappa=1$, was developed in [14]. This approach can be described as a “trajectory viewpoint” on the De Giorgi theory, and it is fully quantitative.
In order to carry out this analysis, one needs to have an a priori local quantitative boundedness result for weak solutions to (1.1), also known as the first De Giorgi lemma, which writes in its simplest form as follows.